An electron and a proton, having the same kinetic energy, enter a region of uniform magnetic field moving perpendicular to the field. Calculate the ratio of the radii of their circular paths in the field.
1
Expert's answer
2019-08-28T09:20:16-0400

According to Newton's second law, we can write

"r=\frac{mv} {Bq} (1)"

The speed is equal to

"v=\sqrt {\frac{2EK } {m} } (2)"

We put (2) in (1)

"\frac{r_p}{r_e}=\sqrt{\frac{m_p}{m_e}} (3)"

where mp=1.67×10-27 kg, me=9.1×10-31 kg

 

Using (3) we got

"\frac{r_p} {r_e}=42.8"


Answer

42.8